### v4

@@ 3.1 Convergence Rate Analysis
    The covariance matrix (22) is not calculable, as so many unknowable variables intervene in its definition. It should consequently be estimated using multiple runs of the MLSA scheme.
    The ES variance factor could notwithstanding be estimated by approximating $\mathbb{V}\mathrm{ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})$ as in (11) with
    $\mathbb{V}\mathrm{ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})\approx\frac{1}{N_{0}}\sum_{k=1}^{N_{0}}\big((X_{h_{0}}^{(k)}-\xi^{h_{0}}_{k-1})^{+}\big)^{2}-\bigg(\frac{1}{N_{0}}\sum_{k=1}^{N_{0}}(X_{h_{0}}^{(k)}-\xi^{h_{0}}_{k-1})^{+}\bigg)^{2},$
@@
    and $\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)$ with
    $\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)\approx\frac{1}{N_{L}}\sum_{k=1}^{N_{L}}\big(\mathds{1}_{X_{h_{L}}^{(k)}>\xi^{h_{L}}_{k-1}}G_{h_{L}}^{(k)}\big)^{2}-\bigg(\frac{1}{N_{L}}\sum_{k=1}^{N_{L}}\mathds{1}_{X_{h_{L}}^{(k)}>\xi^{h_{L}}_{k-1}}G_{h_{L}}^{(k)}\bigg)^{2}.$
@@
    [18, Theorem 2.11] establishes a CLT for biased multilevel schemes under a mean-reverting property [18, Assumption (HMR)] that only concerns the VaR component of our MLSA scheme. Additionally, the martingale array studied in the proof of the aforementioned theorem is different from the one in Step 9 of the proof of Theorem 3.5, Appendix E. Indeed, due to the discontinuity of the gradient $H_{1}$ (6), the summation order of the martingale array had to be swapped to allow invoking the CLT for martingale arrays [27, Corollary 3.1].
@@ 3.2 Complexity Analysis
The global error of the MLSA scheme (17)-(18) decomposes into statistical and bias errors:
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{0}_{\star}$ $=\big(\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)+(\xi^{h_{L}}_{\star}-\xi^{0}_{\star}),$ (23)
$\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{0}_{\star}$ $=\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)+(\chi^{h_{L}}_{\star}-\chi^{0}_{\star}).$ (24)
@@
$L=\bigg\lceil\frac{\ln{(h_{0}/\varepsilon)}}{\ln{M}}\bigg\rceil\sim\frac{\left|\ln{\varepsilon}\right|}{\ln{M}}$ (25)
achieves a global convergence rate in distribution for the MLSA scheme (17)-(18) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. In view of (23)-(24), Lemma A.3(i) and Theorem 3.5 guarantee that the convergence rate of the MLSA scheme is of order $h_{L}$ as $L\uparrow\infty$. Thus, to achieve an error of order $\varepsilon$, we must choose $h_{L}\leq\varepsilon$, hence (25) follows. The remaining complexity computations are standard and are thereby omitted.
@@
MLSA is computationally optimal for $\beta=1$ under the condition $\lambda\gamma_{1}>1$. It runs optimally in $\mathrm{O}(\varepsilon^{-\frac{5}{2}})$ time, an order of magnitude faster than NSA and ANSA that run optimally in $\mathrm{O}(\varepsilon^{-3})$ time.
@@ 4 Averaged Multilevel Stochastic Approximation Algorithm
As shown in Theorem 3.8, the MLSA scheme (17)-(18) is optimal for $\beta=1$ under the non trivial constraint $\lambda\gamma_{1}>1$. The hyperparameter $\gamma_{1}$ must be carefully fine-tuned, adding a significant tuning phase to MLSA’s execution time. To address this limitation, we look into applying the Polyak-Ruppert averaging principle to the multilevel paradigm.
For each level $0\leq\ell\leq L$ and $j\in\{(\ell-1)^{+},\ell\}$, setting $\bar{\xi}^{h_{j},\ell}_{0}=0$, we consider the estimate $\overline{\xi}^{h_{j},\ell}_{N_{\ell}}$ calculated by averaging out the simulations $\{\xi^{h_{j},\ell}_{n},1\leq n\leq N_{\ell}\}$ from (19). The AMLSA estimator of the VaR is defined by
$\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}=\overline{\xi}^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\overline{\xi}^{h_{\ell},\ell}_{N_{\ell}}-\overline{\xi}^{h_{\ell-1},\ell}_{N_{\ell}}$. (26)
@@ 4.1 Convergence Rate Analysis
$N_{\ell}=h_{L}^{-2}\bigg(\sum_{\ell^{\prime}=0}^{L}h_{\ell^{\prime}}^{-\frac{1}{4}}\bigg)h_{\ell}^{\frac{3}{4}},\quad 0\leq\ell\leq L,$ (27)
@@
$\begin{pmatrix}h_{L}^{-1}\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)\\ h_{L}^{-\frac{9}{8}}\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\overline{\Sigma}^{\text{\tiny\rm ML}})\quad\mbox{ as }\quad L\uparrow\infty,$
@@
$\overline{\Sigma}^{\text{\tiny\rm ML}}=\begin{pmatrix}\frac{\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)^{2}(1-M^{-1/4})}&0\\ 0&\frac{h_{0}^{-3/8}(1-M^{-1/4})^{1/2}\mathbb{V}\mathrm{ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})}{(1-\alpha)^{2}}+\frac{h_{0}^{1/4}\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}G)}{(1-\alpha)^{2}M^{1/4}}\end{pmatrix}$, (28)
@@
    The iterations amounts in (27) are determined in Appendix G.2.
@@
    The convergence rate for the VaR is in $\mathrm{O}(h_{L})$, as with MLSA. The convergence rate for the ES is in $\mathrm{O}(h_{L}^{\frac{9}{8}})$ for any $\beta\in\big(\frac{8}{9},1\big)$, without any extra condition on $\gamma_{1}$. This rate matches the optimal convergence rate of the ES MLSA scheme, albeit with the additional constraint $\lambda\gamma_{1}>1$.
@@
    The covariance matrix (28) is not determinable numerically, but it is estimatable using multiple runs of the AMLSA scheme.
@@ 4.2 Complexity Analysis
Recalling the ES global error decomposition (24), we decompose similarly the VaR’s global error into a statistical and a bias error:
$\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{0}_{\star}=\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)+(\xi^{h_{L}}_{\star}-\xi^{0}_{\star}).$ (29)
Theorem $4.$
Let $\varepsilon\in(0,1)$ be some prescribed accuracy. Within the framework of Theorem 4.1, if $h_{0}>\varepsilon$, then, setting $L$ as in (25) achieves a global convergence rate in distribution for the AMLSA scheme (26)-(18) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. Lemma A.3(i) and Theorem 3.5 guarantee that, with regard to (29)-(24), the convergence rate of the AMLSA scheme is of order $h_{L}$ as $L\uparrow\infty$. Therefore, to attain an error of order $\varepsilon$, we must ensure that $h_{L}\leq\varepsilon$, hence (25) follows. The remaining cost computations are standard and are therefore skipped.
Remark 4.6.
@@
NSA (7)-(8) Corollary 1.4 $^{\beta} hh$ $\mathcal{H}\ni h\downarrow 0$, $\beta\in\big(\frac{1}{2},1\big]$,
@@
ANSA (13)-(8) Corollary 2.4 $h$ $h$ $\mathcal{H}\ni h\downarrow 0$, $\beta\in\big(\frac{1}{2},1\big)$
MLSA (19)-(20) Theorem 3.5 $h_{L}$ $h_{L}^{\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)}}$ $L\uparrow\infty$, $\beta\in\big(\frac{1}{2},1\big]$,
@@
AMLSA (26)-(20) Theorem 4.1 $h_{L}$ $h_{L}^{\frac{9}{8}}$ $L\uparrow\infty$, $\beta\in\big(\frac{8}{9},1\big)$
@@
NSA (7)-(8) Theorem 1.8 $\varepsilon^{-\frac{3}{\beta}}$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$
ANSA (13)-(8) Theorem 2.8 $\varepsilon^{-3}$, $\beta\in\big(\frac{1}{2},1\big)$
MLSA (19)-(20) Theorem 3.8 $\varepsilon^{-1-\frac{3}{2\beta}}$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$
AMLSA (26)-(20) Theorem 4.4 $\varepsilon^{-\frac{5}{2}}$, $\beta\in\big(\frac{8}{9},1\big)$
@@ 5 Financial Case Study
The goal of this numerical study is to check and compare empirically the validity of the Corollaries 1.4 and 2.4 and Theorems 3.5 and 4.1. For benchmarking purposes, we consider a numerical setting that allows to retrieve the VaR and ES of a financial loss either analytically, using an unbiased scheme or using a biased one like the ones studied in the previous sections. The code for this numerical case study is available at github.com/azarlouzi/avg_mlsa.
@@
$\mathrm{d}S_{t}=\kappa S_{t}\mathrm{d}t+\sigma\mathrm{d}W_{t},\quad\mbox{ i.e. }\quad S_{t}=S_{0}\mathrm{e}^{\kappa t}+\sigma\int_{0}^{t}\mathrm{e}^{-\kappa(s-t)}\mathrm{d}W_{s},$ (30)
@@ 5.1 Analytical and Simulation Formulas
$\bar{K}=S_{0}\frac{\sum_{i=1}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}\,\mathrm{e}^{\kappa T_{i-1}}}{\sum_{i=1}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}}.$ (31)
Note that $i_{\delta}=1$, so that, by (30)-(31) and the fact that $\pm\int_{0}^{\delta}\mathrm{e}^{-\kappa s}\mathrm{d}W_{s}\sim\mathcal{N}\big(0,\int_{0}^{\delta}\mathrm{e}^{-2\kappa s}\mathrm{d}s\big)$,
$X_{0}\mathrel{{\mathop{=}\limits^{\mathcal{L}}}}\eta Y,\quad\mbox{ where }\quad\eta=\bar{N}\sigma\sqrt{\frac{1-\mathrm{e}^{-2\kappa\delta}}{2\kappa}}\sum_{i=2}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}\,\mathrm{e}^{\kappa T_{i-1}}$ (32)
@@
$\alpha=\mathbb{P}(X_{0}\leq\xi^{0}_{\star})=\mathbb{P}(\eta Y\leq\xi^{0}_{\star}),\quad\mbox{ i.e. }\quad\xi^{0}_{\star}=\eta F^{-1}(\alpha),$ (33)
@@
$\chi^{0}_{\star}=\mathbb{E}[X_{0}|X_{0}\geq\xi^{0}_{\star}]=\frac{\eta}{1-\alpha}\,\mathbb{E}[Y\mathds{1}_{\eta Y\geq\xi^{0}_{\star}}],\quad\mbox{ i.e. }\quad\chi^{0}_{\star}=\frac{\eta}{1-\alpha}f\Big(\frac{\xi^{0}_{\star}}{\eta}\Big),$ (34)

### v5

@@ 3.1 Convergence Rate Analysis
    The covariance matrix (3.8) is not calculable, as so many unknowable variables intervene in its definition. It should consequently be estimated using multiple runs of the MLSA scheme.
    The ES variance factor could notwithstanding be estimated by approximating $\operatorname{\mathbb{V}ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})$ as in (1.8) with
    $\operatorname{\mathbb{V}ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})\approx\frac{1}{N_{0}}\sum_{k=1}^{N_{0}}\big((X_{h_{0}}^{(k)}-\xi^{h_{0}}_{k-1})^{+}\big)^{2}-\bigg(\frac{1}{N_{0}}\sum_{k=1}^{N_{0}}(X_{h_{0}}^{(k)}-\xi^{h_{0}}_{k-1})^{+}\bigg)^{2},$
@@
    and $\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)$ with
    $\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)\approx\frac{1}{N_{L}}\sum_{k=1}^{N_{L}}\big(\mathds{1}_{X_{h_{L}}^{(k)}>\xi^{h_{L}}_{k-1}}G_{h_{L}}^{(k)}\big)^{2}-\bigg(\frac{1}{N_{L}}\sum_{k=1}^{N_{L}}\mathds{1}_{X_{h_{L}}^{(k)}>\xi^{h_{L}}_{k-1}}G_{h_{L}}^{(k)}\bigg)^{2}.$
@@
    [18, Theorem 2.11] establishes a CLT for biased multilevel schemes under a mean-reverting property [18, Assumption (HMR)] that only concerns the VaR component of our MLSA scheme. Additionally, the martingale array studied in the proof of the aforementioned theorem is different from the one in Step 9 of the proof of Theorem 3.5, Appendix E. Indeed, due to the discontinuity of the gradient $H_{1}$ (1.3), the summation order of the martingale array had to be swapped to allow invoking the CLT for martingale arrays [27, Corollary 3.1].
@@ 3.2 Complexity Analysis
The global error of the MLSA scheme (3.3)-(3.4) decomposes into statistical and bias errors:
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{0}_{\star}$ $=\big(\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)+(\xi^{h_{L}}_{\star}-\xi^{0}_{\star}),$ (3.9)
$\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{0}_{\star}$ $=\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)+(\chi^{h_{L}}_{\star}-\chi^{0}_{\star}).$ (3.10)
@@
$L=\bigg\lceil\frac{\ln{(h_{0}/\varepsilon)}}{\ln{M}}\bigg\rceil\sim\frac{\left|\ln{\varepsilon}\right|}{\ln{M}}$ (3.11)
achieves a global convergence rate in distribution for the MLSA scheme (3.3)-(3.4) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. In view of (3.9)-(3.10), Lemma A.3(i) and Theorem 3.5 guarantee that the convergence rate of the MLSA scheme is of order $h_{L}$ as $L\uparrow\infty$. Thus, to achieve an error of order $\varepsilon$, we must choose $h_{L}\leq\varepsilon$, hence (3.11) follows. The remaining complexity computations are standard and are thereby omitted. ∎
@@
MLSA is computationally optimal for $\beta=1$ under the condition $\lambda\gamma_{1}>1$. It runs optimally in $\operatorname{O}(\varepsilon^{-\frac{5}{2}})$ time, an order of magnitude faster than NSA and ANSA that run optimally in $\operatorname{O}(\varepsilon^{-3})$ time.
@@ 4 Averaged Multilevel Stochastic Approximation Algorithm
As shown in Theorem 3.7, the MLSA scheme (3.3)-(3.4) is optimal for $\beta=1$ under the non trivial constraint $\lambda\gamma_{1}>1$. The hyperparameter $\gamma_{1}$ must be carefully fine-tuned, adding a significant tuning phase to MLSA’s execution time. To address this limitation, we look into applying the Polyak-Ruppert averaging principle to the multilevel paradigm.
For each level $0\leq\ell\leq L$ and $j\in\{(\ell-1)^{+},\ell\}$, setting $\bar{\xi}^{h_{j},\ell}_{0}=0$, we consider the estimate $\overline{\xi}^{h_{j},\ell}_{N_{\ell}}$ calculated by averaging out the simulations $\{\xi^{h_{j},\ell}_{n},1\leq n\leq N_{\ell}\}$ from (3.5). The AMLSA estimator of the VaR is defined by
$\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}=\overline{\xi}^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\overline{\xi}^{h_{\ell},\ell}_{N_{\ell}}-\overline{\xi}^{h_{\ell-1},\ell}_{N_{\ell}}$. (4.1)
@@ 4.1 Convergence Rate Analysis
$N_{\ell}=h_{L}^{-2}\bigg(\sum_{\ell^{\prime}=0}^{L}h_{\ell^{\prime}}^{-\frac{1}{4}}\bigg)h_{\ell}^{\frac{3}{4}},\quad 0\leq\ell\leq L,$ (4.2)
@@
$\begin{pmatrix}h_{L}^{-1}\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)\\ h_{L}^{-\frac{9}{8}}\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\overline{\Sigma}^{\text{\tiny\rm ML}})\quad\mbox{ as }\quad L\uparrow\infty,$
@@
$\overline{\Sigma}^{\text{\tiny\rm ML}}=\begin{pmatrix}\frac{\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)^{2}(1-M^{-1/4})}&0\\ 0&\frac{h_{0}^{-3/8}(1-M^{-1/4})^{1/2}\operatorname{\mathbb{V}ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})}{(1-\alpha)^{2}}+\frac{h_{0}^{1/4}\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}G)}{(1-\alpha)^{2}M^{1/4}}\end{pmatrix}$, (4.3)
@@
    The iterations amounts in (4.2) are determined in Appendix G.2.
@@
    The convergence rate for the VaR is in $\operatorname{O}(h_{L})$, as with MLSA. The convergence rate for the ES is in $\operatorname{O}(h_{L}^{\frac{9}{8}})$ for any $\beta\in\big(\frac{8}{9},1\big)$, without any extra condition on $\gamma_{1}$. This rate matches the optimal convergence rate of the ES MLSA scheme, albeit with the additional constraint $\lambda\gamma_{1}>1$.
@@
    The covariance matrix (4.3) is not determinable numerically, but it is estimatable using multiple runs of the AMLSA scheme.
@@ 4.2 Complexity Analysis
Recalling the ES global error decomposition (3.10), we decompose similarly the VaR’s global error into a statistical and a bias error:
$\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{0}_{\star}=\big(\overline{\xi}^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)+(\xi^{h_{L}}_{\star}-\xi^{0}_{\star}).$ (4.4)
Theorem 4.3.
Let $\varepsilon\in(0,1)$ be some prescribed accuracy. Within the framework of Theorem 4.1, if $h_{0}>\varepsilon$, then, setting $L$ as in (3.11) achieves a global convergence rate in distribution for the AMLSA scheme (4.1)-(3.4) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. Lemma A.3(i) and Theorem 3.5 guarantee that, with regard to (4.4)-(3.10), the convergence rate of the AMLSA scheme is of order $h_{L}$ as $L\uparrow\infty$. Therefore, to attain an error of order $\varepsilon$, we must ensure that $h_{L}\leq\varepsilon$, hence (3.11) follows. The remaining cost computations are standard and are therefore skipped. ∎
Remark $4.$
@@
NSA (1.4)-(1.5) Corollary 1.5 $^{\beta} hh$ $\mathcal{H}\ni h\downarrow 0$, $\beta\in\big(\frac{1}{2},1\big]$,
@@
ANSA (2.1)-(1.5) Corollary 2.3 $h$ $h$ $\mathcal{H}\ni h\downarrow 0$, $\beta\in\big(\frac{1}{2},1\big)$
MLSA (3.5)-(3.6) Theorem 3.5 $h_{L}$ $h_{L}^{\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)}}$ $L\uparrow\infty$, $\beta\in\big(\frac{1}{2},1\big]$,
@@
AMLSA (4.1)-(3.6) Theorem 4.1 $h_{L}$ $h_{L}^{\frac{9}{8}}$ $L\uparrow\infty$, $\beta\in\big(\frac{8}{9},1\big)$
@@
NSA (1.4)-(1.5) Theorem 1.8 $\varepsilon^{-\frac{3}{\beta}}$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$
ANSA (2.1)-(1.5) Theorem 2.6 $\varepsilon^{-3}$, $\beta\in\big(\frac{1}{2},1\big)$
MLSA (3.5)-(3.6) Theorem 3.7 $\varepsilon^{-1-\frac{3}{2\beta}}$, $\beta\in\big(\frac{1}{2},1\big]$, with $\lambda\gamma_{1}>1$ if $\beta=1$
AMLSA (4.1)-(3.6) Theorem 4.3 $\varepsilon^{-\frac{5}{2}}$, $\beta\in\big(\frac{8}{9},1\big)$
@@ 5 Financial Case Study
The goal of this numerical study is to check and compare empirically the validity of the Corollaries 1.5 and 2.3 and Theorems 3.5 and 4.1. For benchmarking purposes, we consider a numerical setting that allows to retrieve the VaR and ES of a financial loss either analytically, using an unbiased scheme or using a biased one like the ones studied in the previous sections. The code for this numerical case study is available at github.com/azarlouzi/avg_mlsa.
@@
$\mathrm{d}S_{t}=\kappa S_{t}\mathrm{d}t+\sigma\mathrm{d}W_{t},\quad\mbox{ i.e. }\quad S_{t}=S_{0}\mathrm{e}^{\kappa t}+\sigma\int_{0}^{t}\mathrm{e}^{-\kappa(s-t)}\mathrm{d}W_{s},$ (5.1)
@@ 5.1 Analytical and Simulation Formulas
$\bar{K}=S_{0}\frac{\sum_{i=1}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}\,\mathrm{e}^{\kappa T_{i-1}}}{\sum_{i=1}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}}.$ (5.2)
Note that $i_{\delta}=1$, so that, by (5.1)-(5.2) and the fact that $\pm\int_{0}^{\delta}\mathrm{e}^{-\kappa s}\mathrm{d}W_{s}\sim\mathcal{N}\big(0,\int_{0}^{\delta}\mathrm{e}^{-2\kappa s}\mathrm{d}s\big)$,
$X_{0}\mathrel{{\mathop{=}\limits^{\mathcal{L}}}}\eta Y,\quad\mbox{ where }\quad\eta=\bar{N}\sigma\sqrt{\frac{1-\mathrm{e}^{-2\kappa\delta}}{2\kappa}}\sum_{i=2}^{d}\mathrm{e}^{-rT_{i}}\Delta T_{i}\,\mathrm{e}^{\kappa T_{i-1}}$ (5.3)
@@
$\alpha=\mathbb{P}(X_{0}\leq\xi^{0}_{\star})=\mathbb{P}(\eta Y\leq\xi^{0}_{\star}),\quad\mbox{ i.e. }\quad\xi^{0}_{\star}=\eta F^{-1}(\alpha),$ (5.4)
@@
$\chi^{0}_{\star}=\mathbb{E}[X_{0}|X_{0}\geq\xi^{0}_{\star}]=\frac{\eta}{1-\alpha}\,\mathbb{E}[Y\mathds{1}_{\eta Y\geq\xi^{0}_{\star}}],\quad\mbox{ i.e. }\quad\chi^{0}_{\star}=\frac{\eta}{1-\alpha}f\Big(\frac{\xi^{0}_{\star}}{\eta}\Big),$ (5.5)